# Mixed Strategy Constraints in Continuous Games

### v1

@@ 3.1.2. Properties
The primary appeal of chance constraints comes from the ability to vary $\epsilon$ to soften constraints and remove the undesirable dependence of the solution on arbitrarily poor-quality, but feasible, pure strategies. As such, chance constraints are applicable in “dense” problems with many players and significant mixing, in which many joint strategies have low feasibility due to the high number of player interactions. The existence of a solution is maintained by allowing players to mix with all strategies, not only those with no possibility of infeasibility; as the number of pure strategies grows the possibility of infeasibility disappears. In this way we might expect the general amount of mixing to decrease as $\epsilon$ increases.
On the other hand, $\epsilon$ also has an effect on the solution quality; low settings of it can permit low-feasibility equilibria that are filtered out at higher confidences. In games that have mixed strategy equilibria where the constraint is fully upheld, we might expect the amount of mixing to increase with $\epsilon$.
@@ 3.2. An Alternative: Constraints on Expectation
There is a point of view from which chance constraints appear needlessly verbose, particularly as we develop the machinery needed to apply them to pure strategy optimization. Why is the constraint indicating function $\rho$ necessary at all - why not define the $K$-th element of $Q_{i,j,K}$ directly as $g_{i,j}(s_{K})$, in the same manner we define the elements of the cost tensor $A$?
This is interpretable as constraining the expectation of the feasibility, and in fact, expectation constraints are somewhat justifiable in discrete games. They have a well-defined, intuitive meaning when the pure strategies are limited in number and static. Also, when using penalty methods (under a scalar penalty, perhaps a dynamic one), expectation-constrained games become ordinary tensor games with no irregular behavior, aside from that inherited from $f$ and $g$.
@@
We seek to avoid this behavior when using pure strategy optimization, and we will return to this example to determine if this poor case in fact occurs. First, however, we develop a solver capable of optimizing pure strategies in tensor games with tensor constraints in the following section.
4. Implementing Constrained Pure Strategy Optimization
Thus far, we have discussed tensor games with tensor constraints, where the tensor constraints capture feasible set interactions in the presence of strategic mixing. We now proceed to describe our method of performing pure strategy optimization, i.e. choosing both the pure strategies that result in cost and constraint tensors, as well as the mixing weights over these tensors.
As an aside, the method of pure strategy optimization we suggest here is relatively simple: we append the pure strategies to the decision variables of the tensor game, formulate a mixed complementarity problem using the KKT conditions, and solve. However, note that there are other ways of performing pure strategy optimization; for instance, (Peters et al., 2022) formulate it as a two-level optimization by calculating gradients through an ordinary tensor game. We find the simple approach to be faster for the problems we explore in this work, although other approaches an interesting topic for future research.
4.1. Strategy-Augmented Tensor Games with Tensor Constraints
To model equilibrium seeking in both the strategy space and the weight space, the simplest approach is to include the pure strategies as decision variables of the GNEP. This results in a variation on a tensor game with tensor constraints, with the following form:
[…]
@@ 4.1. Strategy-Augmented Tensor Games with Tensor Constraints
(8) $\nabla_{x_{i}}L_{i}(x,s,\lambda,\gamma)$ $\perp 0\leq x_{i}\leq\infty$
@@
with the Lagrangian function:
(9) $L_{i}(x,s,\lambda,\gamma)=A_{i}(s)\;[\cdot]\;x-\lambda_{i}\sum_{k=1}^{m_{i}}x_{i,k}-\sum_{j\in\mathcal{O}_{i}}\gamma_{i,j}(Q_{i,j}(s)\;[\cdot]\;x).$
Unlike in the finite game setting, the optimization problems (7) are generally non-convex with respect to the decision variables. Therefore a solution to the joint KKT conditions (8) is not necessarily a proper equilibrium point. Nevertheless, for our purposes of exploring ways to model constraints in mixed-strategy continuous games, the KKT approach is satisfactory. The mixed complementarity problem implied by (8) can be solved with an ordinary MCP solver in the typical way. The difficulty of solving this MCP depends on the nonlinearity of $f$, $g$, and $\rho$; under linear functions the problem is multilinear and therefore comparable to the MCPs defined by (5).
@@ 4.2. Iterative Tightening
The approach above only holds under twice differentiable cost, constraint, and constraint indicating functions. While it is reasonable to expect this fact about the cost and constraint functions in spatial games, it does not hold for $\rho$ under default chance constraints, because the indicator 1 is non-differentiable. To circumvent this, we “soften” the indicator into a sigmoid function, so that
(10) $\rho(g(s))=\sigma(\omega g(s))=\frac{1}{1+e^{-\omega g(s)}},$
where $\omega$ sets the extent of softening. We seek to use the highest value of $\omega$ that is numerically feasible, since $\lim_{\omega\rightarrow\infty}\sigma(\omega x)=\text{1}(x\geq 0)$. If $\omega$ is too small, the function $\rho$ does not act as a good approximation to the indicator function, and the resulting tensor constraints will not be representative of proper chance constraints. If $\omega$ is too high, the constraint gradient will be lost numerically and the MCP will generally suffer from poor conditioning. Luckily, near the constraint boundary, this conditioning issue diminishes.
@@
ALGORITHM 1 Constrained Strategy/Weight Equilibrium Computation
@@ 4.3.1. Efficiency
Tensor games in general do not scale well in the number of players, as the cost tensors are exponential in size with $N$. This is a known problem studied in the literature, and there are a variety of methods to ameliorate it, from sampling methods (Gemp et al., 2021) to parallelization (Widger and Grosu, 2009) to neural network approaches (Marris et al., 2022). Fortunately, the addition of constraints does not further affect this complexity, as each constraint tensor is exactly the size of the cost tensors (and are often more sparse); scaling is linear with the number of constraints.
For clarity, in this paper, we assume only proper handling of sparsity in tensors, and otherwise use no efficiency modifications.
@@ 4.3.2. Tooling
We implement Algorithm 1 in Julia (Bezanson et al., 2017), a general-purpose mathematical programming language. We use the PATH solver ( (Dirkse and Ferris, 1995)) to find solutions to the underlying MCPs.
5. Evaluating Constraints under Pure Strategy Optimization
With an adequate solver for seeking equilibria in both continuous pure strategies and their weights under constraints, we now turn to characterizing its applicability towards spatial games with intuitive collision constraints. We primarily consider “pursuit-evasion” games. Pursuit-evader games are a useful benchmark in our considerations as in pure strategy dimension of one or higher, the evader is incentivized to mix in apparently “deceptive” behavior, while pure strategies themselves must be placed strategically as well (Peters et al., 2022).
@@ 5.1.2. Solving Hog-Poacher-Ranger
We deployed an implementation of the solver described in Alg. 1 on one thousand randomized initial solutions of the “hog-poacher-ranger” game for several parameter variations of chance constraints. We noted cases where the game was unable to be solved, as well as the average actual chance of violating the constraint, which generally differs from the constraint as computed during solving. We also note the cost of every player (note that the game is zero-sum between the “hog” and “poacher”) and the time taken to solve. Table 1 summarizes these results. We also include the behavior of expectation constraints as described in Section 3.2
@@
5.1.3. Computation Results: N-Player Hog-Poacher-Ranger Game
To benchmark our solver, we also used an $N$-player variant of this game, in which the first player acts as the hog, the $N-2$ subsequent players act as poachers aiming to minimize expected squared distance to their predecessor, and constrained to maintain a minimum distance from their successor. The final player also aims to minimize the expected squared distance to its predecessor. Because this game is well-defined for $N$ players, we can perform a numerical analysis to evaluate the solution time when solving using Algorithm 1 for various values of $N$ as well as various values of $m$, where $m$ is the number of pure strategies held by each of the players. The results of this experiment are listed in the following table.
$m=2$ 3 4 5
N=2 0.0034±0.0005 0.0088±0.0003 0.0092±0.0004 0.0192±0.0006
3 0.0159±0.0007 0.0531±0.001 0.0521±0.0011 0.2213±0.003
4 0.0484±0.0092 0.0692±0.0008 0.2514±0.0025 0.2376±0.001
5 0.0596±0.0011 0.4768±0.0067 0.3692±0.0043 0.594±0.0018
6 0.0946±0.0022 0.4188±0.0008 0.6166±0.0027 1.3409±0.0024
7 0.0847±0.0015 0.9942±0.0017 1.198±0.006 2.8395±0.0149
Table 2. Solve times for the $N$-player variant of the Hog-Poacher-Ranger game, with $m$ actions for each player. Trials were evaluated over $10$ solves per configuration. Average times are shown in seconds, along with $95\%$ confidence estimates coming from the standard error.
@@ 5.1.4. Parameters: Confidence
Our application of chance constraints uses two types of parameter: $\epsilon_{i,j}$ and $\omega$. We explore the impact they have on the resulting game, considering first $\epsilon_{i,j}$. Figure 2 shows the actual average number of times the constraint was upheld in the “hog-poacher-ranger” game with respect to different settings of $\epsilon_{i,j}$ (that is, the percentage of games in which the poacher “escapes”). We averaged across one hundred games per setting and $\omega_{\textrm{des}}=64$ was used; games in which $\omega<10$ after iterative tightening were considered solver failures. The dotted line corresponds to the expected percentage of “escapes” (the exact constraint). We might expect this to become equal to $\epsilon_{i,j}$, but in fact, this only occurs at $\epsilon\in\{0,0.5,1.0\}$. We see that settings above $\epsilon_{i,j}=0.5$ the solution tends to be overly conservative, while below $\epsilon_{i,j}=0.5$, the solution tends to be loose.
@@ 5.1.5. Parameters: Strictness
Figure 3. Effect $\omega$ on constraint value (minimal distance between “ranger” and “poacher”). Actual constraint at distance of 1.
@@ 5.3. Promotion of mixed strategies
We find that the “hog-poacher-ranger” game typically falls in the latter category: its equilibria are generally perfectly mixed for the “poacher” player, varying from this only at sufficiently low $\epsilon_{i,j}$. This is reasonable, based on the equilibria of this game shown in Figure 1; the true equilibria seem to have equal mixing for the poacher and we should expect to approach that as $\rho$ approaches the indicator 1.
In general, correlating the findings of Figs. 5 and 2, we find that the “hog-poacher-ranger” game is not sensitive to settings of $\epsilon$ given a sufficient $\omega$; all equilibrium strategies occur at perfect poacher mixing. However, this is not always the case, particularly as the number of mixed strategies increases. For instance, in a version of the “hog-poacher-ranger” game where the poacher is allowed to mix over three strategies, at $\epsilon<1$ equilibria exist where the poacher player overlaps the ranger player with probability $1-\epsilon$ and mixes equally over the remaining two strategies. Table 3 shows this relationship between $\epsilon$ and the realized feasibility as computed on a simplified “hog-poacher-ranger“ variant where the “hog” is stationary at (0, 0) and the “poacher” has three pure strategies to mix over.
Confidence ($\epsilon$) 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Realized Feasibility 0.1992 0.4043 0.5901 0.6528 0.7013 0.7536 0.8062 0.8797
Table 3. Relationship between the confidence value $\epsilon$ and the realized (exact) feasibility at equilibrium for a modified “hog-poacher-ranger”. Averages of twenty solves per confidence; $\omega_{\textrm{des}}=64$.
@@ 6. Closing Remarks
We end with some brief commentary about important modeling aspects when applying chance constraints under pure strategy optimization not covered here.
@@ 6.1.1. Fairness
In general, tensor games with tensor constraints can admit multiple solutions. This can occur the same constraint is repeated under different ownership - for instance, $\mathcal{O}_{1}=\mathcal{O}_{2}$, $Q_{1,j}=Q_{2,j}$. This is generally undesirable in problems in physical space (particularly with regards to collision constraints in which, loosely speaking, we expect all players to “push” equally against the constraints). Ideally we seek a fair equilibria (Rabin, 1993).
When constraints appear in the same ownership sets, the formulation used here in fact yields a refinement of generalized Nash equilibria known as the variational equilibrium. The variational equilibrium is a sort of “fair” equilibrium, in that players share dual variables $\gamma_{i,j}$ for any particular constraint (as in (Schiro et al., 2013) and (Nabetani et al., 2011)). Unfortunately, this refinement is not sufficient to ensure a unique solution for tensor games with tensor constraints.
@@ 6.2. Conclusion
We have presented a method for constraining equilibrium problems taking place in both pure strategies and mixing weights using a chance constraint based method. From the point of view of Nash equilibrium users - specifically roboticists, control theorists, and motion planners - the ability to handle feasible set interactions without sacrificing either strategic mixing or pure strategy optimality is critical to modeling problems in physical space. We look forward to applying this work to more complex spatial games and anticipate its utility in overcoming a recurring frustration in modeling agent interactions in game theoretic settings.

### v2

@@ 2.4.1. Properties
The primary appeal of chance constraints comes from the ability to vary $\epsilon$ to soften constraints and remove the dependence of the solution on arbitrarily poor-quality, but feasible, pure strategies. As such, chance constraints are applicable in “dense” problems with many players and significant mixing, in which many joint strategies have low feasibility due to the high number of player interactions. The existence of a solution is maintained by allowing players to mix with all strategies, not only those with no possibility of infeasibility. In this way we might expect the general amount of mixing to decrease as $\epsilon$ increases.
On the other hand, $\epsilon$ also has an effect on the solution quality — low settings of it can permit low-feasibility equilibria that are filtered out at higher confidences. In games that have mixed strategy equilibria where the constraint is fully upheld, we might expect the amount of mixing to increase with $\epsilon$.
@@ 2.5. An Alternative: Constraints on Expectation
There is a point of view from which chance constraints appear needlessly verbose. Why not set $\rho$ to the identity and define the $K$-th element of $Q_{i,j,K}$ directly as $g_{i,j}(s_{K})$, in the same manner we define the elements of the cost tensor $A$? This is interpretable as constraining the expectation of the feasibility. They have a well-defined, intuitive meaning when the pure strategies are limited in number and static. Also, when using penalty methods (under a scalar penalty, perhaps a dynamic one), expectation-constrained games become ordinary tensor games with no irregular behavior, aside from that inherited from $f$ and $g$.
@@
We seek to avoid this behavior when solving continuous games, and we will return to this example to determine if this poor case in fact occurs. First, however, we develop a solver to find Nash equilibria in tensor games with tensor constraints in the following section.
3. Approach
We now proceed to describe our method of solving continuous games — choosing both the pure strategies that result in favorable cost and constraint tensors, as well as the mixing weights over these tensors. This is sometimes referred to as pure strategy optimization, and we use a relatively ordinary approach: we append the pure strategy parameters to the decision variables of the tensor game, formulate a mixed complementarity problem using the KKT conditions, and solve. However, there are other ways of performing pure strategy optimization; for instance, Peters et al. (2022) formulate it as a two-level optimization problem, calculating gradients through an ordinary tensor game. We find this simple approach to be faster for the problems we explore in this work, although other approaches an interesting topic for future research.
3.1. Solver Core
We seek a solution to
$\forall i\in\{1..N\}:\begin{bmatrix}\;\;\min\limits_{x_{i}\in\Delta^{m_{i-1}},s_{i}\in\mathcal{S}_{i}^{m_{i}}}\;\;A_{i}(s)\;[\cdot]\;x\;\;\;\;\textrm{s.t.}\;\;\;\;\\[8.0pt] \;\;\;\;\;\;Q_{i,j}(s)\;[\cdot]\;x\geq\epsilon_{i,j}\;\;\forall j\in\mathcal{O}_{i}\\[8.0pt] \end{bmatrix}$ (7)
where $A_{i}(s)$ and $Q_{i,j}(s)$ are those in Eqn. 4.
@@ 3.1. Solver Core
$\nabla_{x_{i}}L_{i}(x,s,\lambda,\gamma)$ $\perp 0\leq x_{i}\leq\infty$ (8)
@@
with the Lagrangian function
$L_{i}(x,s,\lambda,\gamma)=A_{i}(s)\;[\cdot]\;x-\lambda_{i}\sum_{k=1}^{m_{i}}x_{i,k}-\sum_{j\in\mathcal{O}_{i}}\gamma_{i,j}(Q_{i,j}(s)\;[\cdot]\;x).$ (9)
Unlike in the finite game setting, the optimization problems given by Eqn. 7 are generally non-convex with respect to the decision variables. Therefore a solution to the joint KKT conditions (8) is not necessarily a proper equilibrium point. Nevertheless, for our purposes of exploring ways to model constraints in mixed-strategy continuous games, the KKT approach is satisfactory. The mixed complementarity problem implied by Eqn. 8 is then solved with an ordinary MCP solver. The difficulty of solving this MCP depends on the nonlinearity of $f$, $g$, and $\rho$; under linear functions the problem is multilinear and therefore comparable to the MCPs defined by (5).
@@ 3.2. Iterative Tightening
The approach above only holds under twice differentiable cost, constraint, and constraint indicating functions. While it is reasonable to expect this about the cost and constraint functions in the application areas we are concerned with (robotics and control), it does not hold for $\rho$ under default chance constraints, because the indicator 1 is non-differentiable. To circumvent this, we “soften” the indicator into a sigmoid function, so that
$\rho(g(s))=\sigma(\omega g(s))=\frac{1}{1+e^{-\omega g(s)}},$ (10)
where $\omega$ sets the extent of softening. We seek to use the highest value of $\omega$ that is numerically feasible, since $\lim_{\omega\rightarrow\infty}\sigma(\omega x)=\text{1}(x\geq 0)$. If $\omega$ is too small, the function $\rho$ does not act as a good approximation to the indicator function, and the resulting tensor constraints will not be representative of proper chance constraints. If $\omega$ is too high, the constraint gradient will be lost numerically and the MCP will generally suffer from poor conditioning. Near the constraint boundary, this conditioning issue diminishes and a suitable solution can be found.
@@
ALGORITHM 1 Constrained Strategy/Weight Nash Computation
[…]
Three Nash equilibria of the hog-poacher-ranger game. In the first, the hog mixes 50/50 to the north and south, the ranger places itself in the middle, and the poacher mixes 50/50 to the west and east of the ranger. In the second, the hog is located under the ranger (both with full weight) and the poacher is to either side, 50/50 weight. In the third, a line is formed: 0.5 hog, 0.5 poacher, 1.0 ranger, 0.5 poacher, 0.5 hog.
@@ 3.3.1. Efficiency
Tensor games do not scale well in the number of players as the cost tensors are exponential in size with $N$. This is a known problem, and there are a variety of methods to ameliorate it, from sampling methods Gemp et al. (2021) to parallelization Widger and Grosu (2009) to neural network approaches Marris et al. (2022). The addition of constraints does not further affect this complexity, as each constraint tensor is exactly the size of the cost tensors, so scaling is linear with the number of constraints.
In this paper, we assume only proper handling of sparsity in tensors, and otherwise use no efficiency modifications.
@@ 3.3.2. Tooling
We implement Algorithm 1 in Julia Bezanson et al. (2017), a general-purpose mathematical programming language. We use the PATH solver Dirkse and Ferris (1995) to find solutions to the underlying MCPs.
4. Analysis
We now turn to characterizing the solver’s applicability towards spatial games with intuitive collision constraints. As an illustrative scenario we consider a pursuit-evasion game. Pursuit-evader games are a useful benchmark in our considerations as in pure strategy dimension of one or higher, the evader is incentivized to mix in apparently “deceptive” behavior, while pure strategies themselves must be placed strategically as well Peters et al. (2022).
@@ 4.1.2. Solving Hog-Poacher-Ranger
We deployed an implementation of the solver described in Alg. 1 on one thousand randomized initial solutions of the “hog-poacher-ranger” game for several parameter variations of chance constraints. We noted cases where the game was unable to be solved, as well as the average actual chance of violating the constraint, which generally differs from the constraint as computed during solving. We also note the cost of every player (note that the game is zero-sum between the “hog” and “poacher”) and the time taken to solve. Table 1 summarizes these results. We also include the behavior of expectation constraints as described in Section 2.5.
@@ 4.1.3. Parameters: Confidence
Our application of chance constraints uses two types of parameter: $\epsilon_{i,j}$ and $\omega$. We explore the impact they have on the resulting game, considering first $\epsilon_{i,j}$. Figure 2 shows the average number of times the constraint was actually upheld in the “hog-poacher-ranger” game with respect to different settings of $\epsilon_{i,j}$ (that is, the percentage of games in which the poacher “escapes”). We averaged across one hundred games per setting and $\omega_{\textrm{des}}=64$ was used; games in which $\omega<10$ after iterative tightening were considered solver failures. The dotted line corresponds to the expected percentage of “escapes” (the exact constraint). We see that settings above $\epsilon_{i,j}=0.5$ the solution tends to be overly conservative, while below $\epsilon_{i,j}=0.5$, the solution tends to be loose.
@@
[…]
A sigmoid-shaped line plot of achieved average empirical feasibility for different settings of $\epsilon$. The theoretical feasibility is shown as the dotted line where $\epsilon$ and the realized feasibility are the same. The true, sigmoid-shaped feasibility differs from this most at $\epsilon=0.5$.
@@ 4.1.4. Parameters: Strictness
Figure 3. Effect of $\omega$ on constraint value (minimal distance between “ranger” and “poacher”). Actual constraint at distance of 1.
[…]
Visualization of the true distance between the ranger and the poacher in the HPR game, for different $\omega$ and $\epsilon$. All converge to the correct value as $\omega$ increases. Higher $\epsilon$ values show steeper curves in this convergence.
@@
4.1.5. Computation Results: N-Player Hog-Poacher-Ranger Game
To benchmark our solver, we also used an $N$-player variant of this game, in which the first player acts as the hog, the $N-2$ subsequent players act as poachers aiming to minimize expected squared distance to their predecessor, and constrained to maintain a minimum distance from their successor. The final player also aims to minimize the expected squared distance to its predecessor. Because this game is well-defined for $N$ players, we can perform a numerical analysis to evaluate the solution time when solving using Algorithm 1 for various values of $N$ as well as various values of $m$, where $m$ is the number of pure strategies held by each of the players. The results of this experiment are listed in Table 2.
m=1 m=2 m=3 m=4
N=2 3.4 ± 0.5 8.8 ± 0.3 9.2 ± 0.4 19.2 ± 0.6
N=3 15.9 ± 0.7 53.1 ± 1.0 52.1 ± 1.1 221.3 ± 3.0
N=4 48.4 ± 9.2 69.2 ± 0.8 251.4 ± 2.5 237.6 ± 1.0
N=5 59.6 ± 1.1 476.8 ± 6.7 369.2 ± 4.3 594.0 ± 1.8
N=6 94.6 ± 2.2 418.8 ± 0.8 616.6 ± 2.7 1340.9 ± 2.4
N=7 84.7 ± 1.5 994.2 ± 1.7 1198.0 ± 6.0 2839.5 ± 14.9
Table 2. Timing benchmarks (in milliseconds) on the hog-poacher-ranger game.
@@ 4.2. Outlier Sensitivity of Expectation Constraints
[…]
The “hog” player is depicted in the far corners of the square play area, each corner at 50 percent weight. The “ranger” player sits between them at full weight. The “poacher” player has two pure strategies: one is low-weight and far away from the ranger, and the other intersects the ranger.
@@ 4.3. Promotion of mixed strategies
We find that the “hog-poacher-ranger” game typically falls in the latter category: its equilibria are generally perfectly mixed for the “poacher” player, varying from this only at sufficiently low $\epsilon_{i,j}$.
In general, correlating the findings of Figs. 5 and 2, we find that this game is not sensitive to settings of $\epsilon$ given a sufficient $\omega$. However, this is not necessarily the case for all games. In a version of the “hog-poacher-ranger” game where the poacher is allowed to mix over three strategies, at $\epsilon<1$ equilibria exist where the poacher player overlaps the ranger player with probability $1-\epsilon$ and mixes equally over the remaining two strategies. Table 3 shows this relationship between $\epsilon$ and the realized feasibility on this variant.
Confidence ($\epsilon$) Realized Feasibility
0.2 0.1992
0.3 0.4043
0.4 0.5901
0.5 0.6528
0.6 0.7013
0.7 0.7536
0.8 0.8062
0.9 0.8797
Table 3. Relationship between the confidence value $\epsilon$ and the realized (exact) feasibility at equilibrium for a modified “hog-poacher-ranger”.
@@ 5.1.1. Fairness
In general, tensor games with tensor constraints can admit multiple solutions. This can occur the same constraint is repeated under different ownership — for instance, $\mathcal{O}_{1}=\mathcal{O}_{2}$, $Q_{1,j}=Q_{2,j}$. This is generally undesirable in problems in physical space (particularly with regards to collision constraints in which, loosely speaking, we expect all players to “push” equally against the constraints). Ideally we seek a fair equilibria Rabin (1993).
When constraints appear in the same ownership sets, the formulation used here in fact yields a refinement of generalized Nash equilibria known as the variational equilibrium. The variational equilibrium is a sort of “fair” equilibrium, in that players share dual variables $\gamma_{j}$ for any particular constraint (as in Schiro et al. (2013) and Nabetani et al. (2011)). Unfortunately, this refinement is not sufficient to ensure a unique solution for tensor games with tensor constraints.
@@ 5.2. Conclusion
We have presented a method for constraining equilibrium problems taking place in both pure strategies and mixing weights using a chance constraint based method. From the point of view of Nash equilibrium users — specifically roboticists, control theorists, and motion planners — the ability to handle feasible set interactions without sacrificing either strategic mixing or pure strategy optimality is critical to modeling problems in physical space. We look forward to applying this work to more complex spatial games and anticipate its utility in modeling robot interactions in game theoretic settings.